1B-1. Test the following ODEs for exactness, and find the general solution for those which are exact. a) 3x2y dx + (23 +) dy13-7. Solve y = won the Hint: consider die
Answer #1

O Giren ODES: @ → 3xy dx + (x3 +43) dy=0 compared this ODE with Max+Ndy = 0, we get M = 3ohy, N= x²443 1 for be The necessary(6) (2² 4²) dx + (29²x²) dy=0 - Comparing the ODE with the ODE Max+ Ndy o we get M = x²y² , N=y= 22 e Now , OM = -24, 2N = -2(do 2xy dx-vdy=0 comparing ODE (d) with Max+ Ndy=0 we get M = 284, N2-22 IM IN Now , AM = 2x, ax = -2x so the d. ODE & is not1 * If Mxt Ny 70 & the ODE be homogeneous, then I ntny) is an integrating factor of the canalion Max+ Ndy =or Here, the giveea Gireen - - (3 + x) dy + y dx = 0 - ① comparing ② with Maxt Ndy = 0, we get M=Y, Now, N=-(3 + x) , - So Ran is not exact inHere, 1/2 M ax -- y - y alone. - 11/23 (-1-1) -- 2 = function of the integrating factor is: I Fi = eS ( 2 ) dy - 2:28nco) mul